# SpikySpace: A Spiking State Space Model for Energy-Efficient Time Series Forecasting

### v1

@@ III-B Spike Neuron Model
III-C State Space Model
State Space Models (SSMs) provide a compact, physics-inspired way to model sequence dynamics by evolving a low-dimensional latent state and mapping it to observable outputs. Compared with generic recurrent architectures, SSMs admit exact continuous-to-discrete conversions and an analytic convolutional kernel, which together enable numerically stable long-horizon reasoning and efficient implementations.
@@ III-C State Space Model
Thus, SSMs provide a principled way to model long-range temporal dependencies while maintaining a compact state representation.
IV Method
IV-A Overall Framework
Figure 1: Overall structure of SpikySpace. Left: The structure of SpikingMamba; Right: Selective Scan
Our proposed SpikySpace model is illustrated in Fig. 1. It consists of a residual block, a state space module and a forecasting head. To be specific, it employs a pre-normalization structure, utilizing RMSNorm before the Spiking Mamba unit designed for time series forecasting to stabilize the training dynamics. Thus, it can be represented as:
$\mathbf{y}=\mathbf{x}+\text{SpikingMamba}(\text{RMSNorm}(\mathbf{x}))$ (7)
Within each Spiking Mamba block, we first expands the feature dimension from the number of variance $d_{\text{v}}$ to $2d_{\text{hidden}}$ via an input projection, splitting the output into a data path $\mathbf{x}_{\text{in}}$ and a residual gate path $\mathbf{x}_{\text{res}}$. Here, we denote $d_{\text{v}}$ as the number of variants, using as the input dimension, while $d_{\text{hidden}}$ is the hidden size of the model. In our model, we set the hidden size as a small value to downsize the parameter size. After feature extraction, the data path $\mathbf{x}_{\text{proj}}$ is processed by our spiking SSM block for global dependency capture.
After the spiking SSM blocks, the final feature representation $\mathbf{Z}\in\mathbb{R}^{B\times L\times D_{\text{v}}}$ is permuted and passed through a linear layer as a forecasting head to map the input window length $L_{\text{in}}$ to the required prediction horizon $L_{\text{out}}$.
To align computation with spike events, activations are quantized and converted into equivalent spike trains, and the corresponding matrix-vector product becomes an accumulation over spike indices. This design removes high-cost multiplications with accumulation and lets compute traffic scale with firing rate rather than wall-clock sequence length.
Furthermore, expensive exponentials or divisions in standard activations are replaced by piecewise linear or power-of-two approximations of Softplus and SiLU, which are natively supported on neuromorphic substrates via additions and bit-shifts. Overall, our design provides fixed-latency streaming updates, a constant-size state independent of $L$, and an energy profile that closely tracks the number of synaptic operations.
IV-B State Space Module
The key to SpikySpace’s efficiency is the integration of the Spike Neuron into the core of the SpikingMamba, fundamentally transforming the continuous-time system into a sparse, event-driven mechanism. Our SSM unit is responsible for the global sequence modeling within SpikingMamba.
IV-B1 Local Convolutions with Spiking Activation
Before entering the SSM block, the data path $\mathbf{x}_{\text{in}}$ undergoes depth-wise separable 1D convolution to capture the local temporal feature. Crucially, the non-linear activation following the convolution is replaced by a spike neuron, denoted as $\mathcal{SN}(\cdot)$, which transforms the continuous feature $\mathbf{x}_{\text{in}}$ into a sparse, binary spike train $\mathbf{s}\in\{0,1\}$ over $T$ timesteps:
$\mathbf{s_{\text{in}}}=\mathcal{SN}(\mathbf{x}_{\text{in}}),$ (8)
$\mathbf{s}=\mathcal{SN}(\text{Conv1D}(\mathbf{s}_{\text{in}}),$ (9)
where Conv1D denotes a depthwise convolution applied independently to each latent channel.
IV-B2 Spiking SSM
Then we input $s$ into the SSM block. We use the SSM block to model long-range temporal dependencies through a continuous-time state evolution. Given the input spike train $\mathbf{s}$, it learns a latent dynamical system parameterized by $(\boldsymbol{\Delta},\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$, where $\boldsymbol{\Delta}$ controls the discrete time step and $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ define the system dynamics. Through iterative state updates, the SSM propagates information efficiently over long horizons. Before the selective scan, we derive the step size and dynamic parameters as follows.
After a linear transformation, we expand the input and split it into step-size $\boldsymbol{\Delta}$ and learnable weight matrix $\mathbf{B,C}$. As $\Delta$ is going to be send to a fully connection layer, we also encode it into a spike train. Then we quantize the output into integers to send into our proposed PTSoftplus to get the $\Delta$ we send into the selective scan block.
Thus, the process can be represented as:
$\boldsymbol{\Delta},\mathbf{B},\mathbf{C}$ $=\text{Split}(\mathbf{W}\mathbf{s}+\mathbf{b}),\quad\boldsymbol{\Delta}_{s}=\mathcal{SN}\!\left(\boldsymbol{\Delta}\right),$ (10)
$\tilde{\boldsymbol{\Delta}}$ $=\mathcal{Q}(\mathbf{W}_{\Delta}\boldsymbol{\Delta}_{s}+\mathbf{b}_{\Delta}),$
$\boldsymbol{\Delta}$ $=\mathrm{PTSoftplus}\!\left(\tilde{\boldsymbol{\Delta}}\right),\quad\boldsymbol{\Delta}_{s}=\mathcal{SN}\!\left(\boldsymbol{\Delta}\right),$
where $\mathrm{PTsoftplus}$ is our approximation to replace the complicated activation function $\mathrm{Softplus}$ which will be introduced in Section IV-C and $\mathcal{Q}$ representing a quantization process to ensure the input of $\mathrm{PTsoftplus}$ is an integer.
Meanwhile, the SSM’s diagonalizable dynamics parameters are learned as
$\mathbf{A}\;=\;-\exp(\mathbf{A}_{\log})\in\mathbb{R}^{d_{\text{hidden}}\times n},\qquad\mathbf{D}\in\mathbb{R}^{d_{\text{hidden}}}$ (11)
Here, $\mathbf{A}{\log}$ denotes the logarithm of a trainable parameter $\mathbf{A}$. Since $\mathbf{A}{\log}$ is fixed after training, its exponential counterpart $\exp(\mathbf{A}_{\log})$ can be pre-computed, incurring no extra overhead during inference. Then, $\{\boldsymbol{\Delta},\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}\}$ will be send to the selective scan module.
IV-B3 Spike-Driven Selective Scan Module
With the learned parameters and step size, the SSM then performs a spike-driven selective scan to iteratively update the latent state.
In the original Mamba formulation, the transition matrix is obtained through an exponential operation:
$\overline{\mathbf{A}}_{t}=exp(\boldsymbol{\Delta}_{t}\mathbf{A})$ (12)
To facilitate hardware efficiency, we approximate the exponential with a power-of-two representation and round the input to integers, which allows efficient bit-shifting implementation:
$\overline{\mathbf{A}}_{t}=2^{\lfloor\boldsymbol{\Delta}_{t}\mathbf{A}\rceil}$ (13)
The state vector $\mathbf{h}_{t}$ acts as the memory. To enforce SNN constraints within the sequential scan loop, we introduce an additional spike neuron after the state update. At each timestep $t$, the discrete time dynamics are defined as:
$\overline{\mathbf{B}}_{t}$ $=\boldsymbol{\Delta}_{t}\mathbf{B}_{t},$ (14)
$\mathbf{h}_{t}$ $=\mathcal{SN}(\overline{\mathbf{A}}_{t}\mathbf{h}_{t-1}+\overline{\mathbf{B}}_{t}\mathbf{u}_{t}),$ (15)
where $\mathbf{h}_{t}\in\mathbb{R}^{B\times d_{\text{hidden}}\times n}$ denotes the hidden state which is encoded into a spike train, and $\mathbf{s}_{t}$ is the spiking input at step $t$. The module output $\mathbf{y}_{t}$ is computed using the $\mathbf{C}$ matrix and a direct skip connection $D$:
This recurrent scanning mechanism allows the model to propagate information over long horizons with linear-time complexity $O(L)$, while maintaining spiking sparsity and discrete-time stability. By dynamically adjusting $\boldsymbol{\Delta}_{t}$, the selective scan adaptively modulates the update rate of each latent dimension, thus achieving efficient state evolution under event-driven constraints. The whole procedure of our SpikingMamba block is shown in Algorithm 2.
Algorithm 2 SpikingMamba with Selective Scan
1:Input: Spike train $\mathbf{s}\in\mathbb{R}^{B\times L\times d}$; hidden dimension $d_{\text{hidden}}$; sequence length $L$
2:Require: Linear Projection Weight $\mathbf{W}\in\mathbb{R}^{d\times(r_{B}+r_{C}+r_{\Delta})}$, $\mathbf{W}_{\Delta}\in\mathbb{R}^{r_{\Delta}\times r_{\Delta}}$, Bias $\mathbf{b}\in\mathbb{R}^{r_{B}+r_{C}+r_{\Delta}}$, $\mathbf{b}_{\Delta}\in\mathbb{R}^{r_{\Delta}}$, SSM dynamics $\mathbf{A}_{\log}\in\mathbb{R}^{d_{\text{hidden}}\times n}$,  $\mathbf{D}\in\mathbb{R}^{d_{\text{hidden}}}$
3:Output: Predicted Spike Train $\mathbf{y}=\{\mathbf{y}_{t}\}_{t=1}^{L}$
4:Stage 1: Parameters Preparation
[…]
11:Stage 2: Selective Scan
12:for $t=1$ to $L$ do
13:  $\overline{\mathbf{A}}_{t}\leftarrow 2^{\lfloor\boldsymbol{\Delta}_{t}\mathbf{A}\rceil}$ {Power-of-two Transition}
14:  $\overline{\mathbf{B}}_{t}\leftarrow\boldsymbol{\Delta}_{t}\mathbf{B}_{t}$
15:  $\mathbf{h}_{t}\leftarrow\overline{\mathbf{A}}_{t}\mathbf{h}_{t-1}+\overline{\mathbf{B}}_{t}\mathbf{s}_{t}$
16:  $\mathbf{h}_{t}\leftarrow\mathcal{SN}(\mathbf{h}_{t})$
18:end for
19:return $\mathbf{y}$
IV-B4 Output
After obtaining the output $\mathbf{y}$ from the SSM, we further apply a lightweight residual gating module to enhance nonlinearity and stabilize quantized representations. Specifically, we quantize the residual signal, followed by the proposed PTSiLU activation, which will be detailed introduced in IV-D. The processing of $\mathbf{x}_{\text{res}}$ is represented as:
$\mathbf{\tilde{x}}_{\text{res}}=\text{PTSiLU}(\mathcal{Q}\!\left(\mathbf{x}_{\text{res}}\right))$ (17)
Then the residual mechanism would accumulate the $\mathbf{\tilde{x}}_{\text{res}}$ according to the output spike train $\mathbf{y}$ with as:
$\mathbf{\tilde{y}}=\mathbf{y}\odot\mathbf{\tilde{x}}_{\text{res}},$ (18)
where $\odot$ denotes the element-wise product. The final model output is obtained through a linear projection:
$\mathbf{z}=\mathbf{W}_{\text{o}}\mathbf{\tilde{y}}+\mathbf{b}_{\text{o}}.$ (19)
where $\mathbf{W}_{\text{o}}$ is the weight and $\mathbf{b}_{\text{o}}$ is the bias of the output forecasting head. Note that although the input to the final forecasting head, $\tilde{\mathbf{y}}$, is not explicitly encoded as spike trains, it is derived from the element-wise product between a spike train and the output of the PTSiLU activation. Since PTSiLU operates on integer valued inputs using power-of-two scaling, its output consists of quantized values that can be efficiently processed through simple accumulation, bit-shifting, and look-up table operations.
The above formulation completes our SpikingMamba block. In the following section, we detail our proposed activation approximations, PTSoftplus and PTSiLU, which replace conventional exponential-based functions with power-of-two formulations, enabling bit-shift implementation on neuromorphic hardware.
IV-C Power-of-Two Softplus
The Softplus function is widely used in SSMs such as Mamba, where the non-negativity of state parameters and stable gradient flow are crucial for modeling long-range temporal dynamics. The continuous and differentiable nature of Softplus provides a numerically stable alternative to the hard ReLU. Formally, given an input tensor $\mathbf{x}\in\mathbb{R}^{B\times T\times D}$, where $B$ denotes the batch size, $T$ the sequence length, and $D$ the feature dimension, the Softplus activation is defined as:
$\mathrm{Softplus}(\mathbf{x})=\ln\left(1+e^{\mathbf{x}}\right).$ (20)
Each element of $\mathbf{x}$ is independently transformed by this function, producing an output tensor $\mathbf{y}\in\mathbb{R}^{B\times T\times D}$ with the same shape but strictly positive values.
However, while Softplus offers clear algorithmic advantages, its reliance on exponential and logarithmic operations which is hard to implemented by neuromorphic hardware, making direct applying softplus in SNNs impractical. To address this limitation, we propose Power-of-Two based Softplus (PTSoftplus), a piecewise function designed to preserve the smooth and non-negative characteristics of Softplus while substantially reducing computational complexity. This approximation enables efficient and low-power realization of Softplus-like behavior on SNNs and other event-driven hardware, facilitating the deployment of SSMs in energy-constrained neuromorphic environments.
Our approximate Softplus function, PTSoftplus, is defined as the following piecewise function:
[…]
where the parameters
Smoothness and Closeness Analysis
To ensure the stability of this approximation during gradient descent training, we first prove that $\text{PTSoftplus}(x)$ is also a continuously differentiable function.
@@ IV-C Power-of-Two Softplus
Next, we demonstrate that the proposed PTSoftplus function closely approximates Softplus. Specifically, the functional deviation is bounded by a small constant. The derivatives of PTSoftplus and Softplus also exhibit pointwise closeness, ensuring the approximation does not detrimentally affect the model’s training dynamics.
@@
Define $f(x)={\rm PTSoftplus}(x)-{\rm Softplus}(x)$. Then the supremum norm of $f$ (i.e. $\|f\|_{\infty}$) is the maximum deviation between the function values of PTSoftplus and Softplus. Since both functions are continuously differentiable, the function $f$ is also continuously differentiable. The supremum norm of $f^{\prime}$ (i.e. $\|f^{\prime}\|_{\infty}$) bounds the maximum difference between the first-order derivative of PTSoftplus and Softplus.
As $x\rightarrow$-$\infty$, the value of $f(x)$ approaches 0, and as $x\rightarrow\infty$, the value of $f(x)$ approaches the constant value ${1\over\ln 2}-\log_{2}({1\over\ln 2})$. An analysis of the positivity of $f^{\prime}(x)$ on the entire domain $\mathbb{R}$ can show that the function $f$ is monotonically increasing on $\mathbb{R}$. So the value of $f(x)$ must be between 0 and ${1\over\ln 2}-\log_{2}({1\over\ln 2})$. That is,
@@
Since $f^{\prime}(x)$ is always positive, to find an upper bound on $\|f^{\prime}\|_{\infty}$, it suffices to find the maximum value of $f^{\prime}(x)$. It can be shown that, for $x\geq x_{c}$, $f^{\prime}(x)$ is strictly decreasing. Therefore, the maximum value of $f^{\prime}(x)$ on $x_{c},\infty)$ is
@@
On the other hand, an analysis of the positivity of $f^{\prime\prime}(x)$ on $(-\infty,x_{c})$ can show that $f^{\prime}(x)$ is strictly increasing when $x<x_{c}$. Therefore, we can conclude that the maximum value of $f^{\prime}(x)$ on $\mathbb{R}$ is below 0.371, and consequently
$\|f^{\prime}\|_{\infty}=\sup_{x\in\mathbb{R}}|f^{\prime}(x)|\leq f^{\prime}(x_{c})\leq 0.371,$
thereby completing the proof. ∎
The detailed proof of Lemma 2 is relegated to the supplementary materials. Fig. 2 visually demonstrates the function shapes and the deviations between the two functions and their respective derivatives.
IV-D Power-of-Two SiLU
The SiLU activation function is a derivative of Softplus and is highly effective in models like Mamba due to its smooth, self-gating properties.
Given an input tensor $\mathbf{x}\in\mathbb{R}^{B\times T\times D}$, where $B$ denotes the batch size, $T$ the sequence length, and $D$ the feature dimension, the SiLU activation is defined as:
$\mathrm{SiLU}(\mathbf{x})={\bf x}\cdot{\rm Sigmoid}({\bf x}).$ (22)
Each element of $\mathbf{x}$ is independently transformed by this function, producing an output tensor $\mathbf{y}\in\mathbb{R}^{B\times T\times D}$ with the same shape.
For hardware efficiency, we investigate a simpler piecewise exponential-linear approximation function to SiLU. The proposed activaiton function, PTSiLU, is defined as:
where the parameters
Smoothness and Closeness Analysis
Similar to the PTSoftplus function, to ensure the stability of this approximation during training, we first prove that PTSiLU is also a continuously differentiable function.
@@ IV-D Power-of-Two SiLU
Next, we show that PTSiLU closely approximates SiLU, evidenced by the bounded difference between their function values. Furthermore, the proximity of their first-order derivatives ensures that this approximation does not impair the convergence or stability of the model’s training process.
@@
Define $\bar{f}(x)={\rm PTSiLU}(x)-{\rm SiLU}(x)$. Then the analysis of the deviations between the function values and derivatives of PTSiLU and SiLU reduces to the analysis of $\bar{f}$ and $\bar{f}$$^{\prime}$. Given the complexity of PTSiLU, we employ a strategy that is different from the one used in Lemma 1.
We start from the second-order derivative of $\bar{f}$, namely $\bar{f}$$^{\prime\prime}$. A detailed analysis of the terms in $\bar{f}^{\prime\prime}(x)$ yields the following upper bound on the absolute value of $\bar{f}^{\prime\prime}(x)$:
$\sup_{x\in\mathbb{R}}|\bar{f}^{\prime\prime}(x)|\leq(\ln 2)^{2}\cdot 2^{-\bar{x}_{c}-1}+0.5\leq 1.332,$
which means the rate of change of $\bar{f}^{\prime}(x)$ is bounded by the constant value 1.332.
Then, we employ a computer-aided analysis for $\bar{f}^{\prime}$ and $\bar{f}$. We generate a sequence of points ${\cal X}=\{x_{1},x_{2},\dots,x_{N}\}$ from -10 to 10, where $x_{i+1}-x_{i}=0.001$ for $i=1,\dots,N-1$. Then we compute the value of $\bar{f}^{\prime}(x)$ for each $x\in{\cal X}$. The numerical results show that these values are between -0.178 and 0.261. By the fact that $\|\bar{f}^{\prime\prime}\|_{\infty}\leq 1.332$, we have

### v2

@@ Appendix A Spike Generation
Appendix B State Space Model
@@
Appendix C Proofs and Analyses
C.1 Proof of Lemma 1
@@
C.2 Proof of Lemma 2
@@
Proof.
Define $f(x)={\rm PTSoftplus}(x)-{\rm Softplus}(x)$. To find an upper bound on the maximum deviation between PTSoftplus and Softplus, it suffices to determine the minimum and maximum values of the function $f$:
$f(x)=\begin{cases}2^{x}-\ln(1+e^{x}),&\text{if }x<x_{c},\\ x-\ln(1+e^{x})+{1\over\ln 2}-\log_{2}({1\over\ln 2}),&\text{otherwise}.\end{cases}$
As $x\rightarrow$-$\infty$, the value of $f(x)$ approaches 0, and as $x\rightarrow\infty$, the value of $f(x)$ approaches the constant value ${1\over\ln 2}-\log_{2}({1\over\ln 2})$. From Lemma 1, we know that $\rm PTSoftplus$$^{\prime}$ is continuous, and since $\rm Softplus$$^{\prime}$ is also continuous, we know that $f$ is continuously differentiable. The derivative of $f$ is
$f^{\prime}(x)=\begin{cases}(\ln 2)\cdot 2^{x}-{e^{x}\over 1+e^{x}},&\text{if }x<x_{c},\\ {1\over 1+e^{x}},&\text{otherwise}.\end{cases}$
It is clear that, for $x\geq x_{c}$, the derivative $f^{\prime}(x)$ is strictly positive, so $f(x)$ is monotonically increasing when $x\geq x_{c}$. Next, we show that $(\ln 2)\cdot 2^{x}>{e^{x}\over 1+e^{x}}$ when $x<x_{c}$, and consequently $f(x)$ is monotonic on $\mathbb{R}$. By the fact that $2^{x}=e^{x\ln 2}$, the inequality can be reformulated as
$(\ln 2)\cdot e^{x\ln 2}>{e^{x}\over 1+e^{x}}.$ (20)
Divide both sides by $e^{x\ln 2}$. Then the above becomes
$\ln 2>{e^{x(1-\ln 2)}\over 1+e^{x}}.$
Define $g(x)$ as the function on the right-hand side, where
$g^{\prime}(x)={e^{x(1-\ln 2)}\cdot(1-\ln 2-(\ln 2)\cdot e^{x})\over(1+e^{x})^{2}}.$
Since $(1+e^{x})^{2}$ and $e^{x(1-\ln 2)}$ are always positive, setting the first derivative $g^{\prime}(x)$ to 0 gives
$1-\ln 2-(\ln 2)\cdot e^{x}=0\quad\Rightarrow\quad x=\ln\Big({1-\ln 2\over\ln 2}\Big).$
Denote the critical point $x_{g}=\ln({1-\ln 2\over\ln 2})$. It can be easily checked that $g^{\prime}(x)>0$ when $x<x_{g}$, and $g^{\prime}(x)<0$ when $x>x_{g}$. Therefore, $g$ is a concave function whose maximum value is
$g(x_{g})={e^{x_{g}(1-\ln 2)}\over 1+e^{x_{g}}}\leq 0.540<\ln 2\ (\approx 0.693),$
which validates the inequality (20). Consequently, $f^{\prime}(x)$ is strictly positive on $\mathbb{R}$, and $f$ is a monotonic function. Since $f(x)\rightarrow 0$ as $x\rightarrow$-$\infty$ and $f(x)\rightarrow{1\over\ln 2}-\log_{2}({1\over\ln 2})$ as $x\rightarrow\infty$, we conclude that
@@
Next, we show that the difference between the derivatives of PTSoftplus and Softplus is also bounded – that is, the supremum norm of $f^{\prime}$ is bounded. From the above analyses, we already know that $f^{\prime}(x)$ is always positive on $\mathbb{R}$. It remains to find an upper bound on the value of $f^{\prime}(x)$.
It is clear that for $x\geq x_{c}$, $f^{\prime}(x)$ is strictly decreasing. Therefore, on $x_{c},\infty)$, the maximum value of $f^{\prime}(x)$ is
@@
For $x<x_{c}$, the second-order derivative of $f$ is
$f^{\prime\prime}(x)=(\ln 2)^{2}\cdot 2^{x}-{e^{x}\over(1+e^{x})^{2}}.$
In what follows, we show that the above $f^{\prime\prime}(x)>0$ on $(-\infty,x_{c})$, thus $f^{\prime}(x)$ is strictly increasing. In other words, we want to show
$(\ln 2)^{2}>{e^{x(1-\ln 2)}\over(1+e^{x})^{2}}.$ (21)
Define $h(x)$ to be the function on the right-hand side, whose derivative is
$h^{\prime}(x)=\frac{e^{x(1-\ln 2)}\cdot(1+e^{x})\cdot[(1-\ln 2)(1+e^{x})-2e^{x}]}{(1+e^{x})^{4}}.$
Since $e^{x(1-\ln 2)}\cdot(1+e^{x})$ is strictly positive, setting $h^{\prime}(x)$ to zero gives
$(1-\ln 2)(1+e^{x})-2e^{x}=0\quad\Rightarrow\quad x=\ln\Big({1-\ln 2\over 1+\ln 2}\Big).$
Denote the critical point $x_{h}=\ln({1-\ln 2\over 1+\ln 2})$. It can be easily checked that $h^{\prime}(x)>0$ when $x<x_{h}$, and $h^{\prime}(x)<0$ when $x>x_{h}$. Therefore, $h$ is a concave function whose maximum value is
$h(x_{h})={e^{x_{h}(1-\ln 2)}\over(1+e^{x_{h}})^{2}}\approx 0.424\leq(\ln 2)^{2}\ (\approx 0.480),$
which validates the inequality (21). Consequently, $f^{\prime\prime}(x)$ is strictly positive and $f^{\prime}(x)$ is increasing on $(-\infty,x_{c})$. By the fact that $f^{\prime}(x)$ is continuous, we conclude that
$\|f^{\prime}\|_{\infty}=\sup_{x\in\mathbb{R}}|f^{\prime}(x)|=f^{\prime}(x_{c})\leq 0.371,$
which completes the proof. ∎
C.3 Proof of Lemma 3
@@
C.4 Proof of Lemma 4
@@
Proof.
Define $\bar{f}(x)={\rm PTSiLU}(x)-{\rm SiLU}(x)$. Then the analysis of the deviation between PTSiLU and SiLU reduces to the analysis of $\bar{f}$. We start from the derivatives of $\bar{f}$. The first derivative of $\bar{f}$ is
$\bar{f}^{\prime}(x)=\begin{cases}-(\ln 2)\cdot 2^{x}-{e^{x}\cdot(1+x+e^{x})\over(1+e^{x})^{2}},&\text{if }x<\bar{x}_{c},\\ 1-(\ln 2)\cdot 2^{-x-1}-{e^{x}\cdot(1+x+e^{x})\over(1+e^{x})^{2}},&\text{otherwise}.\end{cases}$
On $(-\infty,\bar{x}_{c})$, the second derivative of $\bar{f}$ is
$\bar{f}^{\prime\prime}(x)=-(\ln 2)^{2}\cdot 2^{x}-{e^{x}(2+x+e^{x}(2-x))\over(1+e^{x})^{3}},$
and on $\bar{x}_{c},\infty)$, the second derivative of $\bar{f}$ is
$\bar{f}^{\prime\prime}(x)=(\ln 2)^{2}\cdot 2^{-x-1}-{e^{x}(2+x+e^{x}(2-x))\over(1+e^{x})^{3}}.$
Denote $\bar{h}^{\prime\prime}(x)$ as the function $e^{x}(2+x+e^{x}(2-x))\over(1+e^{x})^{3}$}, which can be simplified to
$\bar{h}^{\prime\prime}(x)=\underbrace{\frac{2e^{x}}{(1+e^{x})^{2}}}_{A(x)}+\underbrace{\frac{x(e^{x}-e^{2x})}{(1+e^{x})^{3}}}_{B(x)}.$ (22)
Here, the first term $A(x)$ is always positive. For the second term $B(x)$, since $(1+e^{x})^{3}>0$, its sign is determined by the numerator $x(e^{x}-e^{2x})$. When $x<0$, we have $(e^{x}-e^{2x})>0$, so $B(x)<0$; when $x>0$, we have $(e^{x}-e^{2x})<0$, so $B(x)$ is also negative. Therefore, $B(x)$ has a maximum value of 0 at $x=0$. Since $A(x)>0$ and $B(x)\leq 0$, to bound the values of $\bar{h}^{\prime\prime}(x)$, we can find extrema of $A(x)$ and $B(x)$ separately.
The derivative of $A^{\prime}(x)$ is
$A^{\prime}(x)=\frac{2(e^{x}-e^{2x})}{(1+e^{x})^{3}}.$
Setting it to zero gives $x=0$. Moreover, $A^{\prime}(x)>0$ when $x<0$, and $A^{\prime}(x)<0$ when $x>0$. Therefore, $A(x)$ attains its maximum value at $x=0$, where $A(0)=0.5$. For the second term $B(x)$, it can be checked that $B(x)$ is symmetric (i.e. B(x)=B(x)B(x)=B(-x)). Moreover, by the fact that $0<{1-e^{x}\over(1+e^{x})^{2}}<1$ for $x<0$, we have
$B(x)={xe^{x}(1-e^{x})\over(1+e^{x})^{3}}\geq xe^{x}\quad\text{for }x<0.$
The function $xe^{x}$ attains its minimum value of $1\over e$} at x=1x=-1. Therefore, we have $B(x)\geq-{1\over e}\geq-0.368$. Together with (22), we have $\|\bar{h}^{\prime\prime}\|_{\infty}\leq\max\{\sup|A(x)|,\sup|B(x)|\}=0.5$. Therefore, for $\bar{f}^{\prime\prime}$, since $(\ln 2)^{2}\cdot 2^{x}$ is strictly decreasing on $(-\infty,\bar{x}_{c})$, we have
$\sup_{x\in(-\infty,\bar{x}_{c})}|\bar{f}^{\prime\prime}(x)|\leq(\ln 2)^{2}\cdot 2^{\bar{x}_{c}}+\|\bar{h}^{\prime\prime}\|_{\infty}\leq 0.639.$
On the other hand, since $(\ln 2)^{2}\cdot 2^{-x-1}$ is strictly decreasing on $\bar{x}_{c},\infty)$, we have
$\sup_{x\in[\bar{x}_{c},\infty)}|\bar{f}^{\prime\prime}(x)|\leq(\ln 2)^{2}\cdot 2^{-\bar{x}_{c}-1}+\|\bar{h}^{\prime\prime}\|_{\infty}\leq 1.332.$
Combining the above two yields $\|\bar{f}^{\prime\prime}\|_{\infty}\leq 1.332$. In other words, the rate of change of $\bar{f}^{\prime}(x)$ is bounded by 1.332.
Based on the above bound on $\|\bar{f}^{\prime\prime}\|_{\infty}$, we employ a computer aided analysis for $\bar{f}^{\prime}$ and $\bar{f}$. We generate a sequence of points ${\cal X}=\{x_{1},x_{2},\dots,x_{N}\}$ from -10 to 10, where $x_{i+1}-x_{i}=0.001$ for $i=1,\dots,N-1$. Then we compute the value of $\bar{f}^{\prime}(x)$ for each $x\in{\cal X}$. The numerical results show that these values are between -0.178 and 0.261. By the fact that $\|\bar{f}^{\prime\prime}\|\leq 1.332$, we have